FEL Upgrade Project Weekly Brief –July 8-12, 2002




A major milestone was achieved this week with the delivery of the second cryomodule into the FEL vault. Transport of the module was uneventful and the dewar is in approximately the correct location in the beamline. The module now awaits fine positioning by the alignment crew before hookup of the rf waveguides and cooldown to 2K presently scheduled for July 24.


Fred Dylla, Stan Majewski, Andrei Afanasev, Cynthia Keppel of Hampton U., and Alan Todd of AES attended the Workshop on the Role of Nuclear Physics Research Community in Combating Terrorism in Washington, July 11-12. The meeting was sponsored by the DoE Division of Nuclear Physics. We presented information on accelerator applications, nuclear particle detection, THz imaging, and the production of anti-microbial surfaces.

Fred Dylla, Gwyn Williams and George Neil did final dry run presentations in preparation for the JLab Science and Technology Review next week.

George Neil attended a briefing for Dr. Sega, Chief Technologist in the Office of the Secretary of Defense on accelerating funding for the Navy High Power FEL scaleup plan. Also attending were a number of high level military and civilians including Adm. Mathis, and Adm. Cohen. Approval was given to proceed to the next level.

WBS 4 (Injector):

The ball cathode was implanted successfully with the W&M implantation system. The system is now fitted with two roughing pumps, making the plasma significantly more stable. A test run on a sample tube with a kapton sleeve to protect it from arcs will be performed on Monday. The shielded bellows/BPM drawings are in review. The gun stand and chamber were aligned and the SF6 tank was connected so positioning of the high voltage tank could be done.

WBS 6 (RF):

Zone 4 - The SL21 cryomodule was moved to FEL zone 4 this week. The EPICS files are being built and loaded today for this zone. The cooling of this cryomodule is scheduled for July 16 and the control software has to be installed and operating by this time. CHL must be able to control the cryomodule's heaters.

Quarter Klystrons - LCW was plumbed onto the klystron cart for unit 3 this week. Two small leaks were found and fixed on the copper headers. Nearly all interlocks have been wired and tested.

Quarter HVPS - The last of the interlocks are being wired and tested today. Unit 3 will be ready for high voltage next week. The SOP is being circulated comments.

Gun HVPS - The tank for the HVPS was lowered to match the Gun tank, but there is not enough horizontal adjustment. The holes in the tank base plate will be enlarged next week as needed. New resistive material was used this week to make a new Conditioning Resistor. This resistor is drying and will be fired next week.

WBS 8 (Instrumentation):

The BLM front card schematic is done. The drawing has been checked and approved. The pcb layout for the board is well under way and attention is being focused on procurement of parts and device package selection (details). The Fast Raster board check plots have been completed by EECAD and are currently being Q.A.'d for final sign off.

A design review was held for the new Machine Protection System (MPS) this week. The Drive Laser Pulse Controller (DLPC) will be reviewed next week.

All the magnet trim cables (0F through 5F regions) have been pulled to their respective destinations in the vault. Locations and channel assignments for the BLM's have been identified and signal and HV cables have been re-routed and pulled into place.

WBS 9 (Transport):

We are saddened by the sudden death of one of our skilled and dedicated colleagues over the Holiday. Among other duties, Karel Capek was our builder, assembler and installer of diagnostic equipment into the beam line. Without him, there is a huge void in the team of steady handed folks that build the machine and keep it running. We will miss him.


Optical Chicane Dipoles (GW)

• At Magnet Enterprises International (MEI) in Oakland CA. the first GW dipole had its Purcell Gap shims successfully glued down. Variation (with respect to a surface plate) over the entire active area of each pole is of the order of .001 inch. This variation is the combination of two epoxy bonds and two sheets of material plus the flatness of the pole. This is excellent work. In addition to booties and hair nets and clean room, and tack rags, three gluing folks at MEI shaved their heads to make sure no fallen hair got into the glue joints.

• The first unit will now be painted, reassembled and shipped early next week. Three others will follow now that they have the process down.

• Cores for the Second batch of 6 Magnets for UV are almost complete.

• The epoxy from the formulator that is with out flexibilizer was used on coils 14 through 17 and they have no surface welts. This cosmetic problem is solved and the solution will be used at Wang NMR for the other impregnated coils.

Injector Dipoles (DU/DV)

• At Master Machine, The two GU cores were shipped out to Wang NMR for delivery today.

• At Wang NMR the first three GVs were put together. They discovered a few missing machined and hardware parts. The problems were solved by Master Machine shipping the missing items in the above shipment or by Wang purchasing hardware from local suppliers. Wang predicts the GVs will ship by early next week.

Arc 180 Degree Dipoles (GY)

• At Wang NMR, the third GY Coil (of four) is wound.

• GY Coil Potting fixture remains in fabrication.

• Three GG coils were potted with the flexiblized epoxy and look good. Wang NMR will pot the remainder with the non-flexiblized epoxy. Arrangements are being made to exchange epoxy with P. D. George Co., the formulator.

Arc Bend, Reverse Bend Dipoles (GQ, GX)

• Process Equipment Co. (PECo) completed one of the GQ Cores. Gap dimensions were inspected by me on July 11 and passed. Again, extraordinary tolerances (.0003 in.) were held by another manufacturer.

• The shim pieces for the energy change were water jet cut to rough size through 3.5 inches of steel producing thin slices with no distortion. These pre heat-treated slices will be ground to final size using a vacuum chuck.

• At Wang NMR, all eight coils for the GQ Dipoles are wound and are being prepared for potting. GX coil winding has started. We discussed potting fixture construction and methods. They will use vacuum pressure impregnation with the non flexiblized epoxy.


3 inch quad (QX)

• Four more girders were installed on their stands. More girders were assembled, adding to a stockpile that will be installed when personnel are available. The installers have been assigned to the CEBAF during its month-long shut down.

Trim Quad (QT)

• Milhous Control of Virginia finally got two production units to us.

• We are now analyzing the use of the integral GC vertical corrector at several levels of quadrupole operation in anticipation of the change request required for this innovation

Sextupole (SF)

• We are checking the sextupole drawings from DULY Research.

Octupole (OT)

• Work continued on this magnet.

Beam Line and Vacuum

• Master machine started up work on the Arc vacuum Chambers.

• The X Chamber continues fabrication in the JLab Shop.

• Eight girders for the recirculation region in front of the Cryomodule were ordered.

• Design of the four chambers for the optical chicane region is nearing completion

• Design of the regions around the wiggler continues.

• Design of the stands for sextupoles and trim quads is nearing completion.

WBS 11 (Optics):

Progress on the IR upgrade:

Both optical cavity assemblies are being fitted up on the test & assembly stands. This gives us the opportunity to try various combinations of leadscrew/nut combinations to see which has the lowest friction. (Rigorous particulate generation checks will probably wait until we can mount them in the vacuum vessels). We anticipate the mirror mounting and orientation hardware today (7/12) or the first part of next week. We received the final QC on the OCMMS beamline parts. Several small variances (e.g., two planes parallel to 0.003" instead of 0.002") from the drawings were found. We are meeting today to formally review the results and sign off on them. The OCMMS support stands were received. We have firm quotes on the optical breadboards and are awaiting release of funds to order them. The first prototype, compact 50 kW beam dump is brazed. We are getting quotes to have it coated. We received two bids on the optical cavity vacuum vessels. Both were higher than our preliminary estimates, so we are going back to the low bidder to see what drives the price. We anticipate making an award next week.

We have a designer working on the insertable mirrors. Excellent progress is being made, and we will hold a preliminary design review in about a week. We are in the process of developing a LabVIEW program to speed measurements of outgassing rates of optical cavity components.

The prototype backplane cooled mirror holder and a test mirror are mounted in the Mirror Test Stand. We are awaiting water fittings before testing it. The designs for the mirror cassette modifications and the turning cassettes continue to progress.

We have quotes for a new grating that extends the IR range of our spectrometer, and will purchase it. We have been confirming the performance of the ultrafast laser oscillator (amplitude jitter, power & pointing stability, etc) and so far it is exceeding specifications. We have one more measurement (timing jitter) which we'll measure today. The first test of our FROG (FEL pulseshape diagnostic) setup is going well. We received our new IR beam profiler (a Spiricon PyroCAM III) which will improve our capabilities.

Progress on the UV upgrade:

We've received additional optics for the UV optics damage test project. We are setting up the hardware, and are awaiting sign off of the safety documents. We are continuing to survey the literature for the most recent absorption measurements of UV optics. As the IR optical transport system (OTS) design progresses, we are looking ahead at incorporating these design elements into the UV OTS design.

Other activities:

We are pricing new optics for the drive laser OTS.